16. Cahuilla Corporation predicts the following sales in units for the coming four months:

April May June July
Sales in units 320 360 380 320


Each month's ending Finished Goods Inventory in units should be 40% of the next month's sales. March 31 Finished Goods inventory is 128 units. A finished unit requires five pounds of direct material B at a cost of $2.00 per pound. The March 31 Raw Materials Inventory has 280 pounds of direct material B. Each month's ending Raw Materials Inventory should be 30% of the following month's production needs. The budgeted purchases of pounds of direct material B during May should be:

  • 1,822 lbs.

  • 368 lbs.

  • 2,374 lbs.

  • 356 lbs.

  • 1,288 lbs.

22.

Webster Corporation's budgeted sales for February are $315,000. Webster pays sales representatives a commission of 5% of sales dollars. The company pays a sales manager a monthly salary of $3,400 and expects advertising expense of $1,000 per month. Compute the total budgeted selling expenses for February.

  • $20,150.

  • $15,750.

  • $4,400.

  • $19,150.

  • $16,750.

23.

Webster Corporation is preparing a master budget for the first quarter. The company budgets production of 3,080 units in January, 2,800 units in February and 3,740 units in March. Each unit requires 0.6 hours of direct labor. The direct labor rate is $13 per hour. Compute the budgeted direct labor cost for the first quarter budget.

  • $65,520.

  • $75,036.

  • $125,060.

  • $109,200.

  • $57,720.

24.

Flagstaff Company has budgeted production units of 8,200 for July and 8,400 for August. The direct materials requirement per unit is 3 ounces (oz.). The company has determined that it wants to have safety stock of direct materials on hand at the end of each month to complete 20% of the units of budgeted production in the following month. There was 4,920 ounces of direct material in inventory at the start of July. The total cost of direct materials purchases for the July direct materials budget, assuming the materials cost $1.10 per ounce, is:

Multiple Choice

  • $27,192.

  • $32,604.

  • $21,648.

  • $26,928.

  • $27,060.

25.

Flagstaff Company has budgeted production units of 8,700 for July and 8,900 for August. The direct labor requirement per unit is 0.60 hours. Labor is paid at the rate of $19 per hour. The total cost of direct labor budgeted for the month of August is:

  • $99,180.

  • $5,340.

  • $101,460.

  • $5,220.

  • $200,640.
Answer #1
April May June July
Sales 320 360 380 320
Add: Desired Ending Inventory 144 152 128
Total Required 464 512 508
Less: Beginning Inventory 128 144 152
Purchases 336 368 356
Raw material per unit 5 5 5
Raw material required 1680 1840 1780
Add: Desired Ending Inventory 552 534
Less: Beginning Inventory 280 552
Required purchases 1952 1822
Hence, the answer is 1822 lbs
Budgeted Selling Expenses= 315000*5% + 3400 = $19150
Budgeted Labor cost = (3080+2800+3740)*0.6*13 = $75036
Budgeted Labor cost for August = 8900*0.6*19 = $101,460